A contractor agreed to pay a penalty if he failed to remodel a certain building within a specified length of time. The penalties for excess time were P400 for the first day and P100 additional for each succeeding day. If the total penalties was P27,000, how many excess days did he use in completing the job?
I dont have any idea on how to solve this....i hope you can help

Answer by rfer(16322)

27000-400=26600/100=266
add one day for the first days $400 fine +1=267 days